Calculate the volume of 7.0M NaOH required to prepare 300. mL of 0.50M NaOH.
Answer #1

We will solve it by dilution equation M1V1= M2V2

M1= initial concentration = 7 M   V1= initial volume =?

M2= final concentration = 0.50 M   V2= final volume = 300 mL

V1 = M2V2/M1= 0.50*300/7 =21.43 mL